OpenStudy (mathmath333):

For a scholarship, at most n candidates out of 2n + 1 can be selected. If the number of different ways of selection of at least one candidate is 63, the maximum number of candidates that can be selected for the scholarship is:

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{For a scholarship, at most n candidates out of 2n + 1 can be selected.}\hspace{.33em}\\~\\ & \normalsize \text{ If the number of different ways of selection of at least one candidate}\hspace{.33em}\\~\\ & \normalsize \text{ is 63, the maximum number of candidates that can be selected for } \hspace{.33em}\\~\\ & \normalsize \text{ the scholarship is: } \hspace{.33em}\\~\\ & a.)\ 3 \hspace{.33em}\\~\\ & b.)\ 4 \hspace{.33em}\\~\\ & c.)\ 2 \hspace{.33em}\\~\\ & d.)\ 5 \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

This how I understand the question. (2n+1) choose 1 + (2n+1) choose 2 +... +(2n+1) choose n = 63

OpenStudy (anonymous):

solve for n?

OpenStudy (mathmath333):

but how u get the equation i didnt understand

OpenStudy (mathmath333):

yes 3 is the answer

OpenStudy (anonymous):

the number of ways to choose 1 candidate from 2n+1 candidates is (2n+1) choose 1 the number of ways to choose 2 candidates from 2n+1 candidates is (2n+1) choose 2 ... the number of ways to choose n candidates from 2n+1 candidates is (2n+1) choose n and the directions say find the number of ways to choose at least 1, but the maximum is n so add the above

OpenStudy (anonymous):

then set equal to 63 . And you can use the identities to simplify this sum

OpenStudy (anonymous):

once you have solved for n, n is also your maximum number of scholarship candidates

OpenStudy (mathmath333):

is it something like \(2^{x}-1=63\)

OpenStudy (anonymous):

if the sum was from i=1 to 2n+1, then yes , it would be 2^(2n+1) -1 its more complicated

OpenStudy (anonymous):

wolfram gives me a messy expression in terms of Gamma

OpenStudy (anonymous):

you're supposed to do this by hand? correct

OpenStudy (mathmath333):

yes

OpenStudy (anonymous):

there is a symmetry (2n+1) choose 1 = (2n+1) choose ( 2n+1 -1) (2n+1) choose 2 = (2n+1) choose ( 2n+1 -2) etc...

OpenStudy (mathmath333):

the whole thing = 2^(2n+1) -1 right ?

OpenStudy (anonymous):

yes thats the sum from i=1 to 2n+1

OpenStudy (mathmath333):

so 2n+1=6,n=2.5 ?

OpenStudy (anonymous):

its like the pascals triangle, on the way up matches on the way down, relative to the center

OpenStudy (anonymous):

(2^(2n+1) -1 ) / 2 = 63 , seems like a good approximation

OpenStudy (anonymous):

you can use the fact that there is going to be an even number of terms in the pascals triangle, because 2n+1 is odd (and you add one extra value for the zeroth term)

OpenStudy (anonymous):

you can get familiarize with binomials by looking at 7 7 C 0 + 7 C 1 + 7 C 2 + 7 C 3 7 C 4 + 7 C 5 + 7 C 6 + 7 C 7 7 C 3 = 7 C 4 7 C2 = 7 C 5 7 C 1 = 7 C 6 7 C 0 = 7 C 7

OpenStudy (anonymous):

(2n+1) C n = (2n+1) C (n+1) (2n+1) C (n-1) = (2n+1) C (n+2) (2n+1) C (n-2) = (2n+1) C (n+3) ... (2n+1) C 1 = (2n+1) C (2n) notice that what is missing from both is (2n+1) C 0 = (2n+1) C (2n+1)

OpenStudy (anonymous):

Lets sum up the binomial terms from i=0 to i= 2n+1, divide it in half (since we only want half the terms) then get rid of the zeroth term 2^(2n+1)/ 2 - (2n+1) C 0 = 63 this will give you n exactly

OpenStudy (anonymous):

this works because 2n+1 is odd, which produces an even number of terms in the pascal triangle row

OpenStudy (anonymous):

2^(2n+1)/ 2 - (2n+1) C 0 = 63 2^(2n+1) /2 - 1 = 63 2^(2n+1) /2 = 64 2^(2n+1) = 128 n = 3